# Perception-aware Sampling for Scatterplot Visualizations

### v1

@@ 4.1. Perception-aware compression
We create a compressed representation of the dataset $\mathcal{D}$ over the two dimensions relevant to the target visualization. Our compression mechanism splits the visualization canvas in a quad-tree hierarchical partitioning fashion: any cell that fails to meet certain data distribution or perception homogeneity requirements is split into four smaller cells. The partitioning terminates with a collection of non-overlapping bounding boxes of different sizes, such that each box (1) approximates well the distribution of data within that box, and (2) covers areas with roughly similar perception weights. We explain these criteria below.
Data approximation. Our goal is to use the compressed representation to draw samples uniformly at random within each box, instead of accessing the data itself. Therefore, this stopping criterion ensures that indeed, a sample drawn uniformly at random from the box is close enough to the true data points of the box. We use the Chamfer (pseudo)-distance (CD) as our metric. CD is a commonly used metric in the computer vision literature for designing generative neural network models for reconstructing 3D point clouds (Achlioptas et al., 2018; Fan et al., 2017; Wu et al., 2021). It compares the similarity of two unordered sets of points $\mathcal{S}_{1}$, $\mathcal{S}_{2}$ and is defined as follows:
@@
Homogeneity of perception weights. Since we generate points randomly within each box, we need the true data within the box to have similar perception weights. We split a box if the variance of the perception weights of the data within the box is higher than a preset threshold $\sigma$.
Figure 10(a) shows a visual representation of the compression of a dataset for varied values of $\lambda$ and $\sigma$. These thresholds affect the number of boxes in the final compression. Intuitively, more boxes result in lower compression and lower distortion of the final visualization. In this case, we see the lowest compression in the upper left corner and the highest compression in the lowest right corner of Figure 10(a). Figure 10(b) shows the corresponding approximate visualizations for each compression (produced by our ApproPAwS, discussed next). We note that even at the lowest compression (most boxes), the size of the representation is only 1.2 MB; the resulting visualization achieves perceptual similarity on par with PAwS, with a runtime that is about 100x faster (Section 6.2).
@@ 4.2. Approximate PAwS
Given a perception-aware compression, our ApproPAwS algorithm (Algorithm  2) generates a sample with points drawn from the compressed representation, rather than from the actual data. Intuitively, ApproPAwS generates and maintains a small representative dataset $\mathcal{R}$ on the fly, and selects points greedily from $\mathcal{R}$ using the same combined perception and coverage objective as PAwS. The algorithm initializes $\mathcal{R}$ by generating uniformly at random a small, constant number $C$ of points from every box in the compressed representation (lines 57). Using $C>1$ makes the algorithm robust to poor random draws (e.g., a box never getting picked due to a poor representative). When a point is selected for the sample, it gets a replacement from the same box in the representative set (line 13). Even though ApproPAwS does not access the actual data, the construction of the compressed representation ensures that the random points drawn from the boxes conform to the data distribution and perception levels of the original data.
Running Time. The running time of ApproPAwS is $O(k^{2}+k\ C\ |\mathcal{P}|)$, where $C$ is the number of points we draw from each box, and $|\mathcal{P}|$ is the number of boxes in the compression. When we replace a representative, we need to compute the distance of the new point to the current sample ($O(k^{2})$ factor). The second term is similar to that of PAwS, but note that $C\cdot|\mathcal{P}|<<n$, when $C$ is a small constant and $|\mathcal{P}|<<n$. This means that ApproPAwS leads to more gains for large datasets, but when $n$ is small the runtime may actually be worse than PAwS due to the additional factors.
@@ 5. Perceptual Similarity Metrics
In this work, we posit that a sample should not distort the attention focus areas of the original data and express perceptual similarity as the similarity between their saliency maps. This insight allows us to use well-established image-based and distribution-based similarity metrics, which we introduce below.
Metrics for perceptual similarity. Saliency maps are stored and visualized as images; thus, we use a well-established image-based metric as our primary metric: Structural Similarity Index Measure (SSIM) (Wang et al., 2004). We further evaluate four distribution-based metrics from the saliency literature (Bylinskii et al., 2018). These metrics treat a saliency map as a distribution over pixels and measure how accurately a saliency model predicts the attention focus areas of a dataset.
@@
[image-based] is a well-known image similarity metric, frequently used as a loss function in computer vision problems (Zhao et al., 2016; Ledig et al., 2017). The metric assumes two images split into fixed-size windows. The SSIM score is computed as the mean of similarity scores across windows measured over luminance $(\ell)$, contrast $(c)$, and structure $(s)$, i.e., $\textup{SSIM}(\theta_{1},\theta_{2})=\ell(\theta_{1},\theta_{2})^{\alpha}\cdot\c{(}\theta_{1},\theta_{2})^{\beta}\cdot s(\theta_{1},\theta_{2})^{\gamma}$. If $\alpha=\beta=\gamma=1$, all three terms affect the score equally, while each term in the product depends on statistical information like the mean pixel value and variance of pixel values of an image window.
@@
[distribution-based] measures dissimilarity among two probability distributions. We normalize the saliency maps to derive the distributions $p_{\mathcal{D^{\prime}}}$ and $q_{\mathcal{D}}$ and compute JSD as:
@@
Discussion. We use well-established distribution and image-based metrics for measuring the similarity across two saliency maps. Interpreting their behavior and the mechanism based on which they penalize mismatches across saliency maps is a separate research area and is not straightforward (Nilsson and Akenine-Möller, 2020; Bylinskii et al., 2018). As an example, Bylinskii et al. (Bylinskii et al., 2018) find CC is affected symmetrically by false positives and negatives, i.e., indicating an area is salient when it is not and vice versa, while SIM is mostly affected by false negatives. We refer the reader to the in-depth experimental evaluation across different metrics by Bylinskii et al. (Bylinskii et al., 2018) for a better understanding of their behavior, and guidelines on how to use them in practice.
@@ 6. Experimental Evaluation
In this section, we present an extensive evaluation of our perception-aware sampling methods, against five state-of-the-art and baseline approaches, over six datasets with diverse characteristics. Our experiments demonstrate that PAwS robustly outperforms other methods in producing samples that are perceptually similar to the original data, and achieves desirable levels of perceptual similarity at lower sample sizes. It is also orders of magnitude faster than competing state-of-the-art on large datasets (Section 6.1). Moreover, the approximate visualizations of ApproPAwS using the perception-aware compression achieve similar levels of perceptual similarity to PAwS, with 100x gains in runtime performance on our largest dataset of $\sim 3.5$M data points. (Section 6.2).
@@
(a) Epileptic Seizure
(b) Estate Correlation
(c) Estate GIS
(d) MNIST
(e) ACSI
(f) Hidden Correlation
Figure 12. Visualizations of the datasets we use in the experimental evaluation.
@@
is a visualization-aware sampling method for scatterplots (Park et al., 2016). It selects a sample $\mathcal{D^{\prime}}\subseteq\mathcal{D}$ of size $k$ such that for any point $x\in\mathcal{D}$ $\exists\ y\in\mathcal{D^{\prime}}$ that is close enough to $x$. Closeness is controlled through a user-specified parameter $\epsilon$ [^7] Since the data lies in a unit square after normalization, we set $\epsilon=\sqrt{2}/100$ because $\sqrt{2}$ is an upper bound on the diameter value of $\mathcal{D}$. Vas uses the Expand + Shrink (ES) local search algorithm, which starts with a candidate solution of $k$ points and keeps swapping a point at a time until no swap offers improvement. Typically, ES converges to a solution in less than half an hour (Park et al., 2016). Park et al. (Park et al., 2016) discuss a runtime optimization using R-trees, but only recommend it for sample sizes of more than 10K points. For smaller sizes, the overhead of maintaining the R-tree dominates the running time. In this work, we implement the vanilla version of the ES algorithm and refer to the original paper for analyzing improvements in running time. Similar to Max-Min, Vas is effective in preserving data shape and maximizing coverage.
@@
probabilistically over-samples points in sparse areas and under-samples points in denser areas to preserve relative density information (Palmer and Faloutsos, 2000). We use the open-source implementation of Yuan et al. (Yuan et al., 2020) [^8] https://github.com/thu-vis/libsampling/blob/master/sampling/SamplingMethods.py, which employs a KNN neighbors algorithm and samples a point with a probability proportional to its K-farthest neighbor distance. Intuitively, if this distance value is high then the point lies in a sparse area and thus has a higher probability of being selected.
@@
Experimental design. Here, we discuss our methodology in creating the perception-augmented DB, computing perceptual similarity, and selecting sample sizes.
Scatterplot rendering and perception weights. As we discussed in Section 3.1, visual features, such as point size and opacity, affect the saliency information in the visualization. The effect can be significant; e.g., in hidden correlation, the correlation trend is only visible for small marker size and low opacity (Figure 12(f)). Our perception-augmented DBs store aggregate saliency maps over visual configurations varying point size (PS: {2px, 4px, 8px, 16px}) and opacity (OP: {10%, 40%, 70%, 100%}), on a fixed canvas size $14\ \textup{in}\times 12\ \textup{in}$ (i.e., aspect ratio of $1.16$ and resolution $1,085\times 924$). The aggregate saliency maps are projected onto the data to derive saliency weights, which are then augmented with density information. In our implementation, we employ a non-parametric, state-of-the-art, tree-based, density estimator from the scikit-learn library (Bentley, 1975; Pedregosa et al., 2011). [^9] Off-the-shelf density estimators were not scalable for hidden correlation ($\sim 3.5$M), and we approximated the behavior by rendering the original dat using a small point & opacity configuration (i.e., PS=2px, OP=10%).
@@ 6.1. Perceptual quality of sampling methods
In this section, we evaluate perception-aware sampling (PAwS) and the other state-of-the-art methods with respect to their ability to avoid perceptual distortion in the resulting visualizations. We first discuss the behavior of different algorithms, as this is showcased through some qualitative results, and then report on the perceptual similarity scores samples achieve across all metrics. We further evaluate PAwS via a user study, and put results in context with their runtime performance.
Figure 13 showcases samples produced by each method on two of our datasets. Since Max-Min and Vas optimize for coverage, they select points that are evenly distributed across the data space, and can fail to capture trends at small sample sizes. On the other hand, Random and DBS prioritize areas of density, thus failing to capture the overall shape. Blue-noise achieves better balance by adaptively prioritizing the selection of points in sparse and dense areas, but behaves like random sampling on the hidden correlation dataset. PAwS prioritizes points in areas of high perceptual significance without sacrificing coverage, resulting in samples that represent well the overall shape and trends. In Figure 13(b), the PAwS sample identifies the correlation trend, when other methods miss it.
@@
Figure 15. Users’ preferences across all datasets for: (a) the sample that better represents correlation trends, clusters, and other patterns, (b) the sample that better represents outliers and overall data shape, (c) the sample that better represents density variations, and (d) the best overall sample. When users could not pick among samples because of high similarity, they could opt for that option.

### v3

@@ 4.1. Perception-aware compression
We create a compressed representation of the dataset $\mathcal{D}$ over the two dimensions relevant to the target visualization. Our compression mechanism splits the visualization canvas in a quad-tree hierarchical partitioning fashion: any cell that fails to meet certain data distribution or perception homogeneity requirements (discussed below) is split into four smaller cells. The partitioning terminates with a collection of non-overlapping bounding boxes of different sizes, such that (1) data is close to uniformly distributed within each box, and (2) datapoints within each box have roughly similar perception weights. We explain these criteria below.
Data approximation. Our approximation scheme draws uniform random samples from a box in the compressed representation, instead of accessing the actual data. The compression scheme ensures that, for each box, a sample drawn uniformly at random is close enough to true data points. We use the Chamfer (pseudo)-distance (CD), a commonly-used metric in the computer vision literature for designing generative neural network models for reconstructing 3D point clouds (Achlioptas et al., 2018; Fan et al., 2017; Wu et al., 2021). It compares the similarity of two unordered sets of points $\mathcal{S}_{1}$, $\mathcal{S}_{2}$ as follows:
@@
Perception homogeneity. Similarly, because we want the generated datapoints to appropriately represent the true data, the compression also ensures that datapoints within each box have similar perception weights. We split a box if the variance of perception weights within it is higher than a preset threshold $\sigma$.
Figure 10(a) shows a visual representation of the compression of a dataset for varied values of $\lambda$ and $\sigma$. These thresholds affect the number of boxes in the final compression. Intuitively, more boxes result in lower compression and lower distortion of the final visualization. In this case, we see the lowest compression in the upper left corner and the highest compression in the lower right corner of Figure 10(a). Figure 10(b) shows the corresponding approximate visualizations for each compression (produced by ApproPAwS, discussed next). We note that even at the lowest compression (most boxes), the size of the representation is only 1.2 MB; the resulting visualization achieves perceptual similarity on par with PAwS, with a runtime that is about 100x faster (Section 6.2).
@@ 4.2. Approximate PAwS
Given a perception-aware compression, ApproPAwS (Algorithm  2) generates a sample with points drawn from the compressed representation, rather than from the actual data. Intuitively, ApproPAwS operates exactly like PAwS, but instead of performing data selection over the actual data, it generates and maintains a small representative dataset $\mathcal{R}$ on the fly. The algorithm initializes $\mathcal{R}$ by generating uniformly at random a small, constant number $C$ of points from every box in the compressed representation (lines 57). Using $C>1$ makes the algorithm robust to poor random draws (e.g., a box never getting picked due to a poor representative). When a point is added to the sample, a new representative from the same box replaces that point in $\mathcal{R}$ (line 13). Even though ApproPAwS does not access the actual data, the construction of the compressed representation ensures that the random points drawn from the boxes conform to the data distribution and perception levels of the original data.
Running Time. Algorithm  2 has runtime complexity $O(k^{2}+k\ C\ |\mathcal{P}|)$. The first factor is due to computing the distance of a new representative to the current sample at each step. The second factor corresponds to PAwS’s $O(kn)$ complexity, but replaces $n$ with $C\ |\mathcal{P}|$, where $C$ is the number of points we draw from each box (a small constant), and $|\mathcal{P}|$ is the number of boxes in the compression. ApproPAwS has pronounced gains for large datasets ($|\mathcal{P}|<<n$); if $n$ is small its performance may degrade due to the additional factors.
@@ 5. Perceptual Similarity Metrics
A key insight in our work is that a sample should distort saliency as little as possible. We use well-established image- and distribution-based metrics to measure how similar the saliency of a sample is to the saliency of the original data.
Measuring perceptual similarity. Saliency maps are stored and visualized as images; thus, we use a well-established image-based metric as our primary metric: Structural Similarity Index Measure (SSIM) (Wang et al., 2004). We further evaluate four distribution-based metrics from the saliency literature (Bylinskii et al., 2018). These metrics treat a saliency map as a distribution over pixels and measure how accurately a saliency model predicts the attention focus areas of a dataset.
@@
[image-based] is a well-known image similarity metric, frequently used as a loss function in computer vision problems (Zhao et al., 2016; Ledig et al., 2017). With two images split into fixed-size windows, the SSIM score is the mean of similarity scores across windows measured over luminance $(\ell)$, contrast $(c)$, and structure $(s)$, i.e., $\textup{SSIM}(\theta_{1},\theta_{2})=\ell(\theta_{1},\theta_{2})^{\alpha}\cdot\c{(}\theta_{1},\theta_{2})^{\beta}\cdot s(\theta_{1},\theta_{2})^{\gamma}$. If $\alpha=\beta=\gamma=1$, all three terms affect the score equally, while each term in the product depends on statistical information like the mean pixel value and variance of pixel values of an image window.
@@
[distribution-based] measures dissimilarity between two probability distributions. We normalize the saliency maps to derive distributions $p_{\mathcal{D^{\prime}}}$ and $q_{\mathcal{D}}$ and compute:
@@
Discussion. All the above metrics are well-established in the literature. We note that interpreting their behavior and the mechanism based on which they penalize mismatches across saliency maps is a separate research area and is not straightforward (Nilsson and Akenine-Möller, 2020; Bylinskii et al., 2018). For example, CC is affected symmetrically by false positives and negatives (indicating an area is salient when it is not and vice versa), while SIM is mostly affected by false negatives (Bylinskii et al., 2018). We refer the reader to the in-depth experimental evaluation by Bylinskii et al. (Bylinskii et al., 2018) for a better understanding of the metrics’ behavior, and guidelines on how to use them in practice.
@@ 6. Experimental Evaluation
In this section, we present an extensive evaluation, including a user study, of our perception-aware sampling methods, against five state-of-the-art and baseline approaches, over six datasets with diverse characteristics. Our experiments demonstrate that PAwS robustly outperforms other methods in producing samples that are perceptually similar to the original data, and achieves desirable levels of perceptual similarity at lower sample sizes. It is also orders of magnitude faster than competing state-of-the-art on large datasets (Section 6.1). Moreover, the approximate visualizations of ApproPAwS achieve similar levels of perceptual similarity to PAwS, with 100x gains in runtime performance on our largest dataset of $\sim 3.5$M data points. (Section 6.2).
@@
(a) Epileptic Seizure
(b) Estate Correlation
(c) Estate GIS
(d) MNIST
(e) ACSI
(f) Hidden Correlation
Figure 12. Visualizations of datasets used in our evaluation.
@@
is a visualization-aware sampling method for scatterplots (Park et al., 2016). It selects a sample $\mathcal{D^{\prime}}\subseteq\mathcal{D}$ of size $k$ such that for any point $x\in\mathcal{D}$ $\exists\ y\in\mathcal{D^{\prime}}$ that is close enough to $x$. Closeness is controlled through a user-specified parameter $\epsilon$ [^6] Since the data lies in a unit square after normalization, we set $\epsilon=\sqrt{2}/100$ because $\sqrt{2}$ is an upper bound on the diameter value of $\mathcal{D}$. Vas uses the Expand + Shrink (ES) local search algorithm, which starts with a candidate solution of $k$ points and keeps swapping a point at a time until no swap offers improvement. Typically, ES converges to a solution in less than half an hour (Park et al., 2016). Park et al. (Park et al., 2016) discuss a runtime optimization using R-trees, but only recommend it for sample sizes of more than 10K points. For smaller sizes, the overhead of maintaining the R-tree dominates the running time. In this work, we implement the vanilla version of the ES algorithm and refer to the original paper for analyzing improvements in running time.
@@
probabilistically over-samples points in sparse areas and under-samples points in dense areas (Palmer and Faloutsos, 2000). We use an open-source implementation (Yuan et al., 2020) that employs a KNN neighbors algorithm and samples a point with a probability proportional to its K-farthest neighbor distance. [^7] https://github.com/thu-vis/libsampling/blob/master/sampling/SamplingMethods.py A large distance implies that the point is likely in a sparse area and thus has higher probability of being selected.
@@
Experimental design. Here, we discuss our methodology, sample sizes, and other parameter and design choices.
Scatterplot rendering and perception weights. As we discussed in Section 3.1, visual features, such as point size and opacity, affect the saliency of a visualization. The effect can be significant; e.g., in hidden correlation, the correlation trend is only visible for small marker size and low opacity (Figure 12(f)). Our perception-augmented DB stores aggregate saliency maps over visual configurations varying point size (PS: {2px, 4px, 8px, 16px}) and opacity (OP: {10%, 40%, 70%, 100%}), on a fixed canvas size $14\ \textup{in}\times 12\ \textup{in}$ (i.e., aspect ratio of $1.16$ and resolution $1,085\times 924$). The aggregate saliency map is projected onto the data to derive weights, which are then augmented with density information. [^8] Our implementation employs a non-parametric, state-of-the-art, tree-based, density estimator from the scikit-learn library (Bentley, 1975; Pedregosa et al., 2011). Off-the-shelf density estimators were not scalable for hidden correlation ($\sim 3.5$M), and we approximated the behavior by rendering the original data using a small point (2px) & opacity (10%) configuration .
@@ 6.1. Perceptual quality of sampling methods
In this section, we evaluate all methods with respect to their ability to avoid perceptual distortion in the resulting visualizations. We first discuss the behavior of different algorithms, as this is showcased through some qualitative results, and then report on the perceptual similarity scores samples achieve across all metrics. We further evaluate humans’ preferences over samples through a user study, and put results in context with their runtime performance.
Figure 13 showcases samples produced by each method on two of our datasets. Since Max-Min and Vas optimize for coverage, they select points that are evenly distributed across the data space, and can fail to capture trends at small sample sizes. On the other hand, Random and DBS prioritize areas of density, thus failing to capture the overall shape. Blue-noise achieves better balance by adaptively prioritizing the selection of points in sparse and dense areas, but behaves like random sampling on the hidden correlation dataset. PAwS prioritizes points in areas of high perceptual significance without sacrificing coverage, resulting in samples that represent well the overall shape and trends. In Figure 13(b), PAwS is the only method that preserves both the data shape and correlation trend.
@@
Figure 15. In one-to-one matchups against the state-of-the-art, users largely prefer PAwS with respect to (a) representation of correlation trends, clusters, and other patterns, (b) representation of outliers and overall data shape, and (c) representation of density variations. Vas is slightly preferred in the “best overall” category (d), but it is computationally prohibitive. PAwS is preferred overall against all other methods.
@@
Runtime performance. We discuss the runtime performance of the algorithms on the Hidden Correlation dataset ($\sim\!3.5M$), which is the largest one in our evaluation (Figure 14, right). PAwS and Max-Min are the fastest among algorithms with reasonable perceptual efficacy. Their runtimes are similar, because they implement the same algorithm with modified objectives. While Random and DBS are faster, they typically produce samples of poor perceptual quality. The KNN algorithm, which is used by DBS to identify points in sparse areas, dominates its running time, which stays constant across sample sizes. The runtimes of Vas and Blue-noise are significantly longer, by 2–3 orders of magnitude. Vas takes a long time to converge (${\sim}23$ minutes for a sample of 250 points), though we have observed that it produces samples of similar quality after one pass over the data (${\sim}5$ minutes for a sample of 250 points). Figure  14 reports both runtimes (till full convergence and for a single pass). We note that Blue-noise is extremely inefficient (${\sim}3.6$ hours for a sample of 250 points), which aligns with prior analyses (Quadri et al., 2022; Chen et al., 2019).
User study. We conducted a user study to assess how humans perceive PAwS samples against each of the other baselines, across all datasets in our study. We recruited 17 Computer Science students ($\textup{Mean}_{\textup{age}}=27$, $\textup{SD}_{\textup{age}}=3.4$) familiar with statistical concepts, verified through a few simple questions. In each task, participants viewed the full dataset—where they could adjust marker transparency—and compared a PAwS sample against one alternative baseline at a time (a total of five sample pairs). For each sample pair, we asked participants to select the sample that is best at (a) preserving correlation trends, clusters, and other patterns, (b) retaining outliers and overall data shape, and (c) preserving density variations. Finally, we asked participants to select the overall best sample in each pair (Figure 15(d)). For each question, participants could select “sample 1”, “sample 2”, or “no preference”. For each dataset, we chose sample sizes between 0.3 and 16% of the data size (more details in Appendix A).
Figure 15 presents aggregated results across all datasets and participants: i.e., $\sim\!50$ trials per sample pair. For each matchup, we report how often users preferred PAwS, how often they preferred the alternative baseline, and how often they declared no preference. The closest in user preference is Blue-noise, which ends up slightly overtaking PAwS in the “overall best” category, but, as we discussed, Blue-noise is computationally prohibitive. Users show strong preference for PAwS against all other alternative methods. In particular, probabilistic-based methods generally rank below PAwS, but we observed participants only decisively prefer PAwS when those methods fail to capture shape and outliers (as is the case on Hidden Correlation and ACSI). We show a per-dataset analysis of the results in Appendix A.
Key takeaways: PAwS robustly outperforms state-of-the-art sampling methods, with emphasized gains in smaller sample sizes. It is also orders of magnitude faster than the other state-of-the-art. Probabilistic methods have consistently low performance. While Blue-noise often achieves good perceptual similarity, its runtime is prohibitive in large datasets. Diversity-based methods have more stable behavior and achieve similar performance to PAwS for large sample sizes, but Vas is again significantly slower. Our user study further confirms PAwS’ effectiveness, as humans prefer its samples in most cases.
6.2. Performance of approximate visualization
In this section, we evaluate the effectiveness of approximate visualizations generated by ApproPAwS (Algorithm 2), using our perception-aware compression (Section 4.1). Compression splits the canvas area into non-overlapping bounding boxes based on two thresholds, $\lambda$ and $\sigma$, controlling data approximation and homogeneity of perception weights, respectively. These parameters affect the level of compression implicitly: lower thresholds result in more and smaller boxes (lower compression), thus, sample points drawn randomly within a box match better the underlying data and lead to lower distortion (example in Figure 10). We present results at three compression levels: (1) low: $\langle\lambda,\sigma\rangle=\langle 0.001,0.001\rangle$, (2) medium: $\langle\lambda,\sigma\rangle=\langle 0.002,0.001\rangle$, and (3) high: $\langle\lambda,\sigma\rangle=\langle 0.003,0.01\rangle$. For the ACSI dataset, we alter the medium and high levels slightly, to $\langle\lambda,\sigma\rangle=\langle 0.005,0.001\rangle$ and $\langle\lambda,\sigma\rangle=\langle 0.01,0.01\rangle$, respectively. We note that the derived compressions may differ in the number of boxes: e.g., hidden correlation has 9,658, 4,042, and 993 boxes across the three levels (Figure 10(a)), while estate correlation has 3,491, 1,184, and 448 boxes. We give more details and context on these choices in Appendix A. In practice, the thresholds can be tuned by data analysts using a grid search approach to reach the desired level of compression.